OpenStudy (anonymous):

5/8 + 1/6 how did I come to 19/24 simplifying??

OpenStudy (anonymous):

\[\huge \frac{5}{8} +\frac{1}{6}=\frac{5\times 6 + 1 \times 8}{8 \times 6}\] \[\huge =\frac{30 +8}{48}\huge =\frac{38}{48}=\frac{19}{24}\]

OpenStudy (anonymous):

@iriscar3

OpenStudy (anonymous):

is this the simplest form of doing this problem